OpenStudy (anonymous):

Solve the IVP x" - 2tx' + 4x = 0, x(0) = -1, x'(0) = 0. How to start?

OpenStudy (amistre64):

what do you know?

OpenStudy (amistre64):

the initial condtions bring laplace to mind

OpenStudy (amistre64):

variation of parameters comes to mind too

OpenStudy (anonymous):

But... L(2tx')?? Not sure how to do it...

OpenStudy (amistre64):

sL(gf) = L(g'f) + L(gf') + gofo might be useful

OpenStudy (amistre64):

i dont know all the tabled stuff, so i cant just bring something to the mat

OpenStudy (anonymous):

I can't recall I have learnt that formula...

OpenStudy (amistre64):

you know L(y') = sL(y) - yo right?

OpenStudy (anonymous):

Yup!

OpenStudy (amistre64):

by the product rule L(g'f + gf') = sL(gf) - gofo by linearity L(g'f) + L(gf') = sL(gf) - gofo and by addition L(g'f) + L(gf') + gofo = sL(gf)

OpenStudy (amistre64):

not sure how well it plays out, but i know it plays :)

OpenStudy (anonymous):

For variation of parameter, not sure how this question can be solved using that technique since not all coefficients are constants nor do we have the fundamental set of solutions.

OpenStudy (amistre64):

sL(t t) = L(t*1) + L(1*t) + 0,0 sL(t^2) = 2L(t) s^2 L(t^2) = 2s L(t) ----------------------- s L(t*1) = L(1*1) + L(t*0) + 0*1 s L(t) = L(1) s^3 L(t^2) = 2 sL(1) ---------------------- s L(1*1) = L(0) + L(0) + 1 s L(1) = 1 -------------------- s^3 L(t^2) = 2 L(t^2) = 2/s^3

OpenStudy (amistre64):

2s L(t x') = L(1*x') + L(t x'') + 0*x'(0) 2s^2 L(t x') = sL(x') + sL(t x'') 2s^2 L(t x') = s(sL(x)-x(0)) + sL(t x'') --------------------------- sL(t x'') = L(x'') + L(t x''') i see that going in circles, unless we can see a more clever trick to it the table says: L(t x') = -dx/ds

OpenStudy (anonymous):

Just wondering, would it be possible to use series method here? It is odd though, since we are given the initial conditions....

OpenStudy (amistre64):

L(x') = x(s) L(t x') = -x'(s)

OpenStudy (amistre64):

yeah, series is fine

OpenStudy (anonymous):

But then we have to check ordinary point etc etc?!

OpenStudy (amistre64):

dont see why we would have to ....

OpenStudy (anonymous):

We assume the solution being sum from k=0 to infty a_k (t-a)^k, where a = ordinary point?!

OpenStudy (amistre64):

assuming my interpretation of a table is right x" - 2tx' + 4x = 0 s(sL(x)-x(0)) - x'(0) + 4L(x) + 2x'(s) = 0 x(0) = -1, x'(0) = 0 s(sL(x)+1) + 4L(x) + 2x'(s) = 0 s^2 L(x) +s+ 4L(x) + 2x'(s) = 0 (s^2+4) L(x) = - 2x'(s) - s L(x) = -[2x'(s) + s]/[s^2+4]

OpenStudy (anonymous):

Unless I can derive L(t x') = -dx/ds in the exam, I cannot use this result since I have not been taught about this in class...

OpenStudy (amistre64):

well, im saying that it might suggest a cos function?

OpenStudy (amistre64):

well that was a boring solution lol

OpenStudy (anonymous):

How so?!

OpenStudy (amistre64):

i peeked at the wolfs output

OpenStudy (anonymous):

This solution looks so simple, but how to get this... T_T

OpenStudy (amistre64):

t^2 derive to 2t

OpenStudy (anonymous):

Yea, but then?!

OpenStudy (amistre64):

unless youve got an eye for it, lets work a power series

OpenStudy (amistre64):

let x = t^2 + c 4t^2 -2t*2t = 0 2 + 4c = 0 is all thats left c = -1/2 would fit, right?

OpenStudy (amistre64):

oh i forgot the initials .... brain slipped a gear

OpenStudy (anonymous):

How do you know "x = t^2 + c"??

OpenStudy (amistre64):

the wolf, i told you i peeked, and we could most likely guess it if we were brilliant ... but i know im not.

OpenStudy (anonymous):

We don't have wolf in exam lol

OpenStudy (anonymous):

Let \(\phi = \sum_{k=0}^\infty a_k(t)^k\) be the solution. \(\phi' = \sum_{k=1}^\infty ka_k(t)^{k-1}\) \(\phi'' = \sum_{k=2}^\infty k(k-1)a_k(t)^{k-2}\) x'' - 2tx + 4x \[=\sum_{k=2}^\infty k(k-1)a_k(t)^{k-2} - 2t\sum_{k=1}^\infty ka_k(t)^{k-1} + 4\sum_{k=0}^\infty a_k(t)^k\]\[=\sum_{k=2}^\infty k(k-1)a_k(t)^{k-2} - 2\sum_{k=1}^\infty ka_k(t)^{k} + 4\sum_{k=0}^\infty a_k(t)^k\]\[=\sum_{k=0}^\infty (k+2)(k+1)a_k(t)^{k} - 2\sum_{k=1}^\infty ka_k(t)^{k} + 4\sum_{k=0}^\infty a_k(t)^k\]\[=2a_2+4a_0 + \sum_{k=0}^\infty [(k+2)(k+1)a_k(t)^{k} - 2ka_k+4a_k ]t^k\]

OpenStudy (anonymous):

Oh, the last line should be \[=2a_2+4a_0 + \sum_{k=1}^\infty [(k+2)(k+1)a_k(t)^{k} - 2ka_k+4a_k ]t^k\]

OpenStudy (anonymous):

this is basically a short-cut to solving via series expansions: repeatedly differentiate: \[x'' - 2tx' + 4x = 0\\x'''-2tx''+2x'=0\\x^{(4)}-2tx'''=0\\\dots\] now consider at \(x=0\) we have \(x=-1, x'=0\); it's easy to solve to see \(x''=4\) and then from there \(x'''=0\). any furhter derivatives are then trivially going to be such that \(x^{(n)}=0\) for \(n>4\), so for our analytic solution about \(x=0\) we've determined it is at most a quadratic polynomial \(c_0+c_1 x+c_2 x^2\). it's then a matter of solving \(c_0,c_1,c_2\) using \(x=-1,x'=0,x''=4\) at \(x=0\)

OpenStudy (anonymous):

x(0) = -1 => a0 = -1

OpenStudy (anonymous):

and x'(0) = 0 => a1 = 0

OpenStudy (anonymous):

\[4a_0 + 2a_2 = 0\] => a_2 = 2

OpenStudy (anonymous):

Then we have, for the first three terms, -1 + 0(t) + 2(t^2)

OpenStudy (anonymous):

So, now, we have the answer from wolf: 2t^2-1, but then, how do we know the rest of the term in the series are all 0 :S

OpenStudy (anonymous):

@RolyPoly \[[(k+2)(k+1) - 2k + 4]a_k=0\]the only way to guarantee this expression is \(0\) for all integers \(k\) is for \(a_k=0\)

OpenStudy (anonymous):

and the expression msut be \(0\) because it's the coefficient of \(t^k\) in the combined power series expression for \(x''-2tx'+4x\) and our original equation states \(x''-2tx'+4x=0\)

OpenStudy (anonymous):

Oh, right!!! One more question, what do you men when you say "analytic solution", in particular, the word "analytic"?

OpenStudy (anonymous):

Also, do we need to show that a is an ordinary point?!

OpenStudy (thomas5267):

The Laplace Transform of t.f(t). http://math.stackexchange.com/questions/366802/laplace-transform-of-tft